# DReSS: Data-driven Regularized Structured Streamlining for Large Language Models

### v2

@@ 3.2 Regularization
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be efficiently solved using backpropagation (BP) algorithm (Rumelhart et al., 1986). In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example. According to Proposition 3.2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
$\min_{W}\quad\mathcal{L}(W,X)+\lambda||RW_{up}^{i}||_{1}+\lambda||W_{down}^{i}R||_{1}$ (6)
$\min_{W,Y_{1},Y_{2}}$ $\quad\mathcal{L}(W,X)+\mathbf{1}^{T}Y_{1}\mathbf{1}+\mathbf{1}^{T}Y_{2}\mathbf{1}$ (7)
s.t. $\quad-Y_{1}\leq RW_{up}^{i}\leq Y_{1},$
$\quad-Y_{2}\leq W_{down}^{i}R\leq Y_{2},$
$\quad Y_{1}\geq 0,\quad Y_{2}\geq 0.$
During the forward and backward propagation processes, due to the application of regularization on certain specific channels in LLMs, the values within these rows or columns are significantly reduced. Intuitively, this reduction could potentially transfer important information and decrease the impact of these parameters on the model’s performance.
@@ 3.3 Pruning
After completing the regularization, we prune the rows and columns directly using the pseudo-index selection matrix $R$. For clarity, we define $S=I-R$ and the pruning for the $i$th layer attention block is then performed as:
${W_{Q}^{i}}^{\prime}$ $=SW_{Q}^{i},\quad{W_{K}^{i}}^{\prime}=SW_{K}^{i}$ (8)
${W_{V}^{i}}^{\prime}$ $=SW_{V}^{i},\quad{W_{o}^{i}}^{\prime}=W_{o}^{i}S$
While for the FFN block, it follows:
${W_{up}^{i}}^{\prime}=SW_{up}^{i},\quad{W_{down}^{i}}^{\prime}=W_{down}^{i}S$ (9)
@@
${W_{emb}}^{\prime}$ $=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S$ (10)
${w_{N}^{i}}^{\prime}$ $=S{w_{N}^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}$
The entire pruning process is straightforward and simple.
@@ 3.4 Optional RFT
After pruning, we perform RFT on the model using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al., 2021). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still maintains competitive performance in terms of both perplexity and zero-shot accuracy.
Table 1: Performance comparison of different pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy is reported on five zero-shot tasks: PIQA, WinoGrande, HellaSwag, ARC-e, and ARC-c, as well as the average accuracy. The best result is highlighted in bold, and the second-best in underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm.
Model Method Pruning Ratio PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24 LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44 SliceGPT 25% 7.08 69.21 65.35 52.40 53.70 31.66 54.46 SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68 DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47 DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16 OPT-2.7B Dense 0% 12.46 74.81 61.01 60.58 54.42 31.14 56.39 LLM Surgeon 25% 12.60 68.22 58.06 46.25 51.52 27.31 50.27 SliceGPT 25% 14.56 65.29 57.22 47.85 49.79 27.99 49.63 SLEB 25% 16.73 63.28 55.37 43.09 45.62 24.21 46.31 DReSS-$\ell_{2}$ 25% 12.52 69.12 58.98 49.10 50.62 28.15 51.20 DReSS-$\ell_{1}$ 25% 12.48 68.89 58.33 48.62 51.67 27.55 51.01 OPT-6.7B Dense 0% 10.85 76.39 65.19 67.16 60.14 34.64 60.70 LLM Surgeon 25% 11.33 72.12 61.08 62.36 54.19 30.52 56.05 SliceGPT 25% 11.90 70.35 60.62 58.15 52.78 29.52 54.28 SLEB 25% 13.80 71.28 60.96 59.26 51.54 29.83 54.57 DReSS-$\ell_{2}$ 25% 11.06 74.26 63.12 61.26 56.64 31.77 57.41 DReSS-$\ell_{1}$ 25% 11.74 73.86 64.28 60.45 56.36 32.07 57.40 OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79 LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96 SliceGPT 25% 10.94 73.67 64.25 63.28 60.52 34.64 59.27 SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50 DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47 DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07 LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00 LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42 SliceGPT 25% 7.55 66.87 63.38 54.16 58.46 34.56 55.48 SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49 DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54 DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20 LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76 LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61 SliceGPT 25% 6.30 68.55 67.48 58.10 62.50 37.88 58.90 SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70 DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84 DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Figure 3: The ratio of the sum of absolute parameter values after regularization to that before regularization for LLaMA2-7B. For example, regularization is applied to the last 25% of rows in $W_{Q}$, and the absolute sum of these rows after regularization is divided by the sum of the corresponding rows before regularization to obtain the Preserved Ratios. This is applied similarly to other matrices such as $W_{K}$, $W_{V}$, $W_{o}$, $W_{gate}$, $W_{up}$, $W_{down}$, and for layer activations $X$, the sum of the regularized columns is compared to the original sum.
@@ 4 Experiments
This section introduces experimental setup (4.1) and analyzes the effectiveness of DReSS from the following aspects: performance comparison (4.2), acceleration (4.3), robustness under different pruning ratios (4.4), dependency on different datasets (4.5), low overhead (4.6), and ablation study (4.7).
@@ 4.1 Experimental Setup
Implementation: DReSS and baseline methods are implemented on the PyTorch framework (Paszke et al., 2019) , utilizing the Hugging Face Transformers library (Wolf, 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al., 2024a) to evaluate the pruned models. More details are shown in Appendix B.
Datasets: To evaluate the performance of DReSS, we conducted experiments on generation and zero-shot tasks. For generation task, following prior work(Ashkboos et al., 2024), we evaluate the model’s perplexity on WikiText-2. For zero-shot task evaluation, the benchmarks consists of PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), HellaSwag (Zellers et al., 2019), ARC-e and ARC-c (Clark et al., 2018). To demonstrate that DReSS is not specifically dependent on the regularization data, we use a randomly selected subset of data from Alpaca (Taori et al., 2023), WikiText-2 (Merity et al., 2016), PTB (Marcus et al., 1993), and C4 (Raffel et al., 2020) for calibration before pruning and for RFT after pruning.
Models: The models pruned using DReSS include the LLaMA models (LLaMA2-7B, LLaMA2-13B) (Touvron et al., 2023), OPT models (OPT-2.7B, OPT-6.7B, OPT-13B) (Zhang et al., 2022), and Phi-2 (Javaheripi et al., 2023).
Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al., 2023), SliceGPT (Ashkboos et al., 2024), and SLEB (Song et al., 2024). Among them, LLM Surgeon and SliceGPT are channel-wise, SLEB is layer-wise. To ensure a fair comparison, in the subsequent experiments, we ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB before pruning, and that the data used for RFT in each method is also the same. We also compared with powerful unstructured pruning methods like Mangniude (Han et al., 2015), Wanda (Sun et al., 2023), and SparseGPT (Frantar & Alistarh, 2023) and the results are in Appendix C.
Evaluation Metrics: The performance on generation task is measured by perplexity, a well-established and robust metric (Yao et al., 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al., 2024). The acceleration effects are represented by throughput and latency (Song et al., 2024).
@@ 4.2 Performance Comparison
To ensure fairness, we randomly selected 1,000 samples from WikiText-2 for each method with a sequence length of 2048. The ratio of calibartion data and the RFT data was set to 3:1, and the pruning ratio was 25%.
As shown in Table 1, DReSS demonstrates superior performance across both generation and zero-shot tasks achieving lowest perplexity and highest average accuracy across all models. For instance, on LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. Notably, on OPT-13B, compared to the dense model, DReSS experiences only a 1% drop in average accuracy. The performance difference between the $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In Figure 3, we present the ratio of the absolute values of model parameters and activation values after and before regularization of the selected portions. The results show that, in most intermediate layers, the parameters and activation values after regularization are reduced to 30% of their original values, which indicates that the amount of information contained in this part of the parameter space has decreased. We also present the changes occurring in the unregularized part in Appendix D. The results show that this part has increased compared to before, indicating that it now accommodates more important information. Overall, regularization transfers important information hidden in the pruned portions to the remaining parts.
@@ 4.3 Acceleration Effectiveness
Language processing in LLMs comprises two primary stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We separately analyze the speedup achieved in each stage. Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single NVIDIA A100 GPU. Following the methodology of previous work (Song et al., 2024), the token generation test scenario involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
@@
At a pruning ratio of 50% on OPT-13B, DReSS delivers a 35% improvement in throughput and a 30% reduction in latency compared to the dense model. These results highlight the superior efficiency of DReSS in accelerating model inference.
Model Method PR PPL($\downarrow$) Tokens/s TI Latency(ms) Speedup OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00× DReSS 25% 10.38 1194 1.16× 319.42 1.21× DReSS 50% 12.15 1389 1.35× 274.11 1.41× LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00× DReSS 25% 5.12 1215 1.14× 330.8 1.20× DReSS 50% 7.59 1407 1.32× 285.5 1.39×
4.4 Robustness to Varying Pruning Ratios
@@
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B on WikiText-2 under different methods are shown in Figure 4. DReSS significantly outperforms other methods across various pruning ratios. At a pruning ratio of 20%, DReSS performs similarly to the dense model. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to achieve high sparsity levels and unlocking significant potential for model acceleration. A more detailed exploration of pruning ratios using DReSS is provided in the Appendix E.
@@
Figure 4: Perplexity of LLaMA-7B pruned by various approaches under different pruning ratios on WikiText-2.
4.5 Dependency on Datasets
Since DReSS relies on data-driven regularization, we investigate its potential dataset dependence. We evaluated perplexity for four methods on WikiText-2, using calibration and RFT data selected from Alpaca, WikiText-2, PTB, and C4. To ensure fairness, we randomly selected 1,000 samples from each dataset, with other settings consistent with Section 4.2. As shown in Figure 5, DReSS consistently outperforms the other methods across datasets, demonstrating its robustness and scalability. More details are shown in Appendix F.
Figure 5: Comparison of perplexity on different calibration datasets at a pruning ratio of 25% on LLaMA2-7B.

### v3

@@ 3.2 Regularization
We formalize the minimization of the overall loss as an optimization problem. When $\ell_{2}$-norm is used, the problem can be directly solved using backpropagation (BP) algorithm (Rumelhart, Hinton, and Williams 1986). In contrast, when the $\ell_{1}$-norm is employed, as the regularization losses introduced by various components in LLMs are similar, we consider only ~$\tilde{\mathcal{L}}_{\text{FFN}}^{i}$ as an example, the complete objective function transformation is in Appendix C. According to Proposition 2, the optimization problem in Equation 6 can be equivalently transformed into a constrained formulation in Equation 7, enabling its solution by BP algorithm.
During forward and backward propagation, regularization on certain channels in LLMs significantly reduces the values in those rows or columns. Intuitively, this value reduction could decrease the impact of these parameters on the model’s performance thus potentially transfer important information to the unregularized parts.
$\min_{\mathbf{W}}\hskip 5.0pt\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda||\mathbf{RW_{up}^{i}}||_{1}+\lambda||\mathbf{W_{down}^{i}R}||_{1}$ (6)
$\min_{\mathbf{W,Y_{1},Y_{2}}}$ $\quad\mathcal{L}(\mathbf{W},\mathbf{X})+\lambda(\mathbf{1}^{T}\mathbf{Y_{1}}\mathbf{1}+\mathbf{1}^{T}\mathbf{Y_{2}}\mathbf{1})$ (7)
s.t. $\quad-\mathbf{Y_{1}}\leq\mathbf{RW_{up}^{i}}\leq\mathbf{Y_{1}},$
$\quad-\mathbf{Y_{2}}\leq\mathbf{W_{down}^{i}R}\leq\mathbf{Y_{2}},$
$\quad\mathbf{Y_{1}}\geq 0,\quad\mathbf{Y_{2}}\geq 0.$
@@ 3.3 Pruning
According to Figure 2, we prune the regularized rows and columns using the pseudo-index selection matrix $\mathbf{R}$. For clarity, we define $\mathbf{S=I-R}$. The pruning of the $i$th layer attention block and FFN block are performed as:
$\mathbf{{W_{Q}^{i}}^{\prime}}$ $\mathbf{=SW_{Q}^{i},}\quad$ $\mathbf{{W_{K}^{i}}^{\prime}}$ $\mathbf{=SW_{K}^{i},}$ (8)
$\mathbf{{W_{V}^{i}}^{\prime}}$ $\mathbf{=SW_{V}^{i},}\quad$ $\mathbf{{W_{o}^{i}}^{\prime}}$ $\mathbf{=W_{o}^{i}S,}$
$\mathbf{{W_{up}^{i}}^{\prime}}$ $\mathbf{=SW_{up}^{i},}\quad$ $\mathbf{{W_{down}^{i}}^{\prime}}$ $\mathbf{=W_{down}^{i}S}$
@@
$\mathbf{{W_{emb}}^{\prime}=W_{emb}S,\quad{W_{pos}}^{\prime}=W_{pos}S,}$ (9)
$\mathbf{{\alpha^{i}}^{\prime}=S{\alpha^{i}},\quad{\beta^{i}}^{\prime}=S{\beta^{i}},\quad{W_{lm}}^{\prime}=SW_{lm}}$
The pruning process is straightforward and simple.
@@ 3.4 Optional RFT
After pruning, we perform RFT using a small subset of the data selected in Section 3.1, leveraging LoRA (Hu et al. 2022). This step is optional, as experimental results in section 4.7 show that DReSS without RFT still achieves competitive performance in perplexity and zero-shot accuracy.
@@ 4 Experiments
Model Method PR PPL ($\downarrow$) PIQA WinoGrande HellaSwag ARC-e ARC-c Avg_Acc
Phi-2 Dense 0% 5.28 79.11 75.77 73.83 78.32 54.18 72.24
LLM Surgeon 25% 7.26 67.28 63.25 54.24 51.62 35.85 54.44
SliceGPT 25% 7.06 69.32 65.39 52.57 53.78 31.89 54.59
SLEB 25% 7.82 67.94 62.79 49.80 48.55 29.34 51.68
DReSS-$\ell_{2}$ 25% 6.25 68.52 66.71 56.73 52.78 37.63 56.47
DReSS-$\ell_{1}$ 25% 6.28 68.67 65.32 57.16 52.39 37.28 56.16
LLaMA2-7B Dense 0% 5.47 79.11 69.06 75.99 74.58 46.25 69.00
LLM Surgeon 25% 7.38 70.59 65.87 58.66 63.65 38.33 59.42
SliceGPT 25% 7.49 68.15 64.13 56.22 55.39 34.74 55.73
SLEB 25% 10.24 63.25 62.36 53.77 55.82 32.24 53.49
DReSS-$\ell_{2}$ 25% 5.86 73.18 66.49 61.73 65.42 40.86 61.54
DReSS-$\ell_{1}$ 25% 5.81 73.42 65.73 61.92 65.26 39.68 61.20
LLaMA3-8B Dense 0% 5.76 85.56 77.94 79.27 78.84 56.49 75.62
LLM Surgeon 25% 7.62 76.34 70.18 71.46 69.65 49.22 67.37
SliceGPT 25% 8.14 74.37 67.81 69.56 69.83 45.53 65.42
SLEB 25% 11.37 71.68 62.96 66.37 64.61 43.28 61.78
DReSS-$\ell_{2}$ 25% 6.09 78.29 71.17 73.28 72.64 53.62 69.80
DReSS-$\ell_{1}$ 25% 6.12 78.86 70.28 73.85 72.37 53.02 69.68
OPT-13B Dense 0% 10.12 76.82 64.80 69.81 61.87 35.67 61.79
LLM Surgeon 25% 11.02 74.18 64.33 65.37 60.96 34.98 59.96
SliceGPT 25% 10.90 73.72 64.28 63.33 60.59 34.66 59.32
SLEB 25% 12.02 72.62 63.96 62.79 59.21 34.12 58.50
DReSS-$\ell_{2}$ 25% 10.38 74.06 64.57 66.82 61.56 35.33 60.47
DReSS-$\ell_{1}$ 25% 10.56 73.65 64.38 66.46 61.15 34.74 60.07
LLaMA2-13B Dense 0% 4.88 80.47 72.22 79.39 77.48 49.23 71.76
LLM Surgeon 25% 5.75 77.75 69.62 74.31 72.83 43.52 67.61
SliceGPT 25% 6.16 69.69 68.45 63.73 63.46 39.90 61.05
SLEB 25% 7.39 66.93 65.87 55.48 60.06 35.14 56.70
DReSS-$\ell_{2}$ 25% 5.12 76.14 71.22 76.51 73.45 46.87 68.84
DReSS-$\ell_{1}$ 25% 5.16 77.24 70.61 76.75 73.84 45.46 68.78
Table 1: Performance comparison of different pruning methods. ‘PPL’ refers to the perplexity on WikiText-2. The accuracy (%) is reported on five zero-shot benchmarks. The best result is highlighted in bold, and the second-best is underlined. DReSS-$\ell_{2}$ denotes the use of the $\ell_{2}$-norm, while DReSS-$\ell_{1}$ denotes the use of the $\ell_{1}$-norm. The results below are all obtained after RFT.
@@ 4.1 Experimental Setup
Implementation: All methods are implemented in PyTorch (Paszke et al. 2019), using the Hugging Face Transformers library (Wolf 2019). All experiments are conducted on 80GB NVIDIA A100 GPUs. For fairness, we use llm-eval-harness (Gao et al. 2024a) to evaluate.
Datasets: For generation task, we evaluate the model’s perplexity on WikiText-2 (Ashkboos et al. 2024). For zero-shot task, the benchmarks are PIQA (Bisk et al. 2020), WinoGrande (Sakaguchi et al. 2021), HellaSwag (Zellers et al. 2019), ARC-e and ARC-c (Clark et al. 2018). In Appendix H we use data from Alpaca (Taori et al. 2023), WikiText-2 (Merity et al. 2016), PTB (Marcus, Santorini, and Marcinkiewicz 1993), and C4 (Raffel et al. 2020).
Models: We use models commonly adopted in the pruning domain including the LLaMA models (LLaMA2-7B, LLaMA3-8B, LLaMA2-13B) (Touvron et al. 2023; Grattafiori et al. 2024), OPT model (OPT-13B) (Zhang et al. 2022), and Phi-2 (Javaheripi et al. 2023).
Baselines: We evaluate DReSS against competitive structured pruning methods: LLM Surgeon (van der Ouderaa et al. 2024), SliceGPT (Ashkboos et al. 2024), and SLEB (Song et al. 2024).
Evaluation Metrics: The performance on generation task is measured by perplexity(Yao et al. 2022), while zero-shot tasks performance is evaluated using accuracy (Dong et al. 2024). The acceleration effects are represented by throughput and latency (Song et al. 2024).
Model Method PR PPL($\downarrow$) Tokens/s($\uparrow$) TI($\uparrow$) Latency($\downarrow$) Speedup($\uparrow$)
OPT-13B Dense 0% 10.12 1029 1.00× 386.5 1.00×
DReSS 25% 10.38 1194 1.16× 319.42 1.21×
DReSS 50% 13.85 1389 1.35× 274.11 1.41×
LLaMA2-13B Dense 0% 4.88 1066 1.00× 396.9 1.00×
DReSS 25% 5.12 1215 1.14× 330.8 1.20×
DReSS 50% 7.59 1407 1.32× 285.5 1.39×
@@ 4.2 Performance Comparison
For fairness, we randomly selected 1,000 samples from WikiText-2 training set with sequence length of 2048 for each method. We ensure that the data used for regularization in DReSS is consistent with the calibration data used by LLM Surgeon, SliceGPT, and SLEB, and the data used for RFT in each method is also the same. The ratio of calibartion and RFT data was set to 3:1, the reason is discussed in Section 4.7. All methods are implemented under this setup. The pruning ratio was 25%. The last 25% of the values in the diagonal matrix $\mathbf{R}$ are set to 1, which means pruning the last 25% of the rows or columns of the parameter matrix and we discussed the choice of $\mathbf{R}$ in Section 4.4.
As shown in Table 1, DReSS achieves superior performance in both generation and zero-shot tasks. On LLaMA2-7B, its perplexity is 20% lower than the second-best method, LLM Surgeon. On OPT-13B, compared to the dense model, DReSS only drops 1% in average accuracy. The performance difference between $\ell_{2}$-norm and $\ell_{1}$-norm is tiny. In the following, DReSS refers to the $\ell_{2}$-norm version.
@@ 4.3 Acceleration Effectiveness
Language processing in LLMs comprises two primary stages: prompt processing, which is compute-bound, and token generation, which is memory-bound. We separately analyze the speedup achieved in each stage. Table 2 presents the throughput and latency results for OPT-13B and LLaMA2-13B, evaluated using a single 80GB NVIDIA A100 GPU. Following the methodology of previous work (Song et al. 2024), the token generation test involves generating sentences of 128 tokens with a batch size of 64, whereas for prompt processing, latency is measured by processing an input sequence of 2048 tokens.
@@
At a pruning ratio of 50% on OPT-13B, DReSS delivers a 35% improvement in throughput and a 30% reduction in latency compared to the dense model. These results highlight the superior efficiency of DReSS in acceleration.
Cases (1) (2) (3)
PPL($\downarrow$) 5.86 5.89 5.87
Avg_Acc(%) 61.54 61.39 61.46
Table 3: The impact of pruning different channels on LLaMA2-7B. ‘PPL’ is the perplexity on WikiText-2. Avg_Acc is on five zero-shot benchmarks.
4.4 The Impact of Pruning Different Channels
Matrix $\mathbf{R}$ is used to select the channels on which regularization is applied (followed by pruning). We choose three options of $\mathbf{R}$ under 25% sparsity: (1) remove the matrix’s last 25% rows or columns, (2) remove the first 25% rows or columns , (3) divide the matrix into 5 segments, removing each segment’s 5% of the last rows or columns. The results on LLaMA2-7B are shown in Table 3. The performance differences among (1), (2), and (3) are minimal, demonstrating that DReSS is insensitive to pruning different channels.
4.5 Robustness to Different Pruning Ratios
@@
Keeping all other settings consistent with Section 4.2, we extend the pruning ratio from 20% to 60%. The perplexity of LLaMA2-7B under different methods are shown in Figure 3. DReSS significantly outperforms other methods across various pruning ratios. When the pruning ratio is up to 60%, SLEB collapses, while DReSS maintains relatively low perplexity compared to other pruning methods. This demonstrates that DReSS maintains robust performance even under extreme pruning ratios, enabling structured pruning to unlock significant potential for model acceleration. More detailed results are in Appendix G.
@@
Figure 3: Perplexity of LLaMA2-7B pruned by various approaches under different pruning ratios on WikiText-2.
Figure 4: Perplexity of DReSS and other methods on WikiText-2 under different amount of data.
Figure 5: The ratio of the sum of absolute parameter values after regularization to that before regularization on LLaMA2-7B. For example, the absolute sum of the last 25% of rows in $\mathbf{W_{Q}}$ after regularization is divided by the sum of the corresponding rows before regularization to obtain the preserved ratios. This is applied similarly to other matrices $\mathbf{W_{K},W_{V},W_{o},W_{gate},W_{up},W_{down}}$, and layer activations.
Cases N&N&N R&N&N R&N&R N&N&R R&P&N R&P&R N&P&N N&P&R
PPL($\downarrow$) 5.47 5.06 4.86 5.06 5.97 5.63 22.38 11.45
Avg_Acc(%) 69.00 70.39 71.23 70.39 59.56 62.35 47.25 52.73
Table 4: The model performance under different regularization settings. ‘PPL’ is the perplexity on WikiText-2. Avg_Acc is on five zero-shot benchmarks. Pruning ration is 25% and the LLM is LLaMA2-7B. We didn’t perform RFT after pruning.